# Hessian stability and convergence rates for entropic and Sinkhorn potentials via semiconcavity

### v1

@@ 3.2. Quantitative stability estimates of Hessian
Let us consider once again the function $\theta_{s}\coloneqq\psi^{\nu}_{s}-\psi^{\mu}_{s}$ introduced in Section 2 and the forward process $(Y^{\theta}_{s})_{s\in[0,T]}$ defined as $Y^{\theta}_{s}\coloneqq\nabla\theta_{s}(X^{\psi^{\mu},\rho}_{s})$ where $(X^{\psi^{\mu},\rho})_{s\in[0,T]}$ is the Schrödinger bridge (1.3) (from $\rho$ to $\mu$) and recall that
@@
$\mathrm{d}Z^{\theta}_{s}=$ $\,\Big{[}2\,\mathrm{sym}(Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s}))-(Z^{\theta}_{s})^{2}+\nabla^{3}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}\Big{]}\,\mathrm{d}s+\nabla^{3}\theta_{s}(X^{\psi^{\mu},\rho}_{s})\mathrm{d}B_{s}$
$=$ $\,\Big{[}2\,\mathrm{sym}(Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\mu}(X^{\psi^{\mu},\rho}_{s}))+\nabla^{3}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}\Big{]}\,\mathrm{d}s+\nabla^{3}\theta_{s}(X^{\psi^{\mu},\rho}_{s})\mathrm{d}B_{s}\,,$
where for any matrix $M$ the symbol $\mathrm{sym}(M)\coloneqq(M+M^{\intercal})/2$ denotes its symmetrized version and where for any $h\in\{\psi^{\mu}_{s},\,\theta_{s}\}$ and for any vector $v\in\mathbb{R}^{d}$ we have defined the product $\nabla^{3}h\,v$ as the matrix
$(\nabla^{3}h\,v)_{ij}:=\langle\nabla(\partial_{i}\partial_{j}h),\,v\rangle\,.$
@@
22 22 $\frac{Z^{\theta}_{s}\cdot\mathrm{sym}(Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s}))}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}$
$\,=\frac{1}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\biggl{(}\operatorname{Tr}(Z^{\theta}_{s}\cdot Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s}))+\operatorname{Tr}(Z^{\theta}_{s}\cdot\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Z^{\theta}_{s})\biggr{)}$
$\,=\frac{2}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\,\operatorname{Tr}(Z^{\theta}_{s}\cdot\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Z^{\theta}_{s})\geq\frac{2\,\lambda(\psi^{\nu}_{s})\,\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}$
$\,=2\,\lambda(\psi^{\nu}_{s})\,r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})-2\,\lambda(\psi^{\nu}_{s})\,\varepsilon\geq 2\,\lambda(\psi^{\nu}_{s})\,\|Z^{\theta}_{s}\|_{\mathrm{HS}}-2\,\lambda(\psi^{\nu}_{s})\,\varepsilon\,.$
@@
$\frac{Z^{\theta}_{s}\cdot\nabla^{3}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}=\frac{\sum_{ijk}Z^{\theta,ij}_{s}\partial_{ijk}\psi^{\nu}_{s}(X^{\psi^{\mu},\rho}_{s})Y^{\theta,k}_{s}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}$
$\qquad\qquad\geq-\|\nabla^{3}\psi^{\nu}_{s}(X^{\psi^{\mu}}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\,\frac{\|Z^{\theta}_{s}\|_{\mathrm{HS}}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\geq-\|\nabla^{3}\psi^{\nu}_{s}(X^{\psi^{\mu}}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\,.$
@@
$\mathrm{d}r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})\geq\biggl{(}-\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}+2\,\lambda(\psi^{\nu}_{s})\,\|Z^{\theta}_{s}\|_{\mathrm{HS}}-\|\nabla^{3}\psi^{\nu}_{s}(X^{\psi^{\mu}}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\biggr{)}\mathrm{d}s$
$-2\,\lambda(\psi^{\nu}_{s})\,\varepsilon\mathrm{d}s+\frac{Z^{\theta}_{s}\cdot\nabla^{3}\theta_{s}(X^{\psi^{\mu},\rho}_{s})}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\mathrm{d}B_{s}\,.$
@@
$\mathbb{E}\|Z^{\theta}_{0}\|_{\mathrm{HS}}\leq\mathbb{E}\|Z^{\theta}_{t}\|_{\mathrm{HS}}+\int_{0}^{t}\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}\mathrm{d}s-2\int_{0}^{t}\lambda(\psi^{\nu}_{s})\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}\mathrm{d}s$
$+\int_{0}^{t}\mathbb{E}\|\nabla^{3}\psi^{\nu}_{t}(X^{\psi^{\mu},\rho}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\mathrm{d}s\,.$
@@
where the negative part of a real number $a$ is defined as $a^{-}\coloneqq\max\{-a,0\}$. In conclusion, by applying Jensen’s inequality we conclude that
$+\int_{0}^{{\tau_{\ell}}}\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}\mathrm{d}s+\int_{0}^{{\tau_{\ell}}}\mathbb{E}\|\nabla^{3}\psi^{\nu}_{t}(X^{\psi^{\mu},\rho}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\mathrm{d}s\,,$
which ends our proof.
@@
which generalizes the constant $C^{\psi^{\nu}}_{\delta}$ considered in Theorem 3.1. By repeating the same argument employed in Theorem 2.1 when proving the upper bound (2.7) for $\mathbb{E}[|Y^{\theta}_{T}|^{2}]=\mathbb{E}[|Y^{\eta}_{0}|^{2}]$, we can prove the following generalization.
[…]
@@
By reasoning as in the proof of Theorem 2.1, from Itô’s formula and (2.2), for all $s\leq\delta T$ we have
which combined with Grönwall’s lemma gives for any $s\geq\delta^{\prime}T$
$\mathbb{E}[|Y^{\theta}_{\delta^{\prime}T}|^{2}]\,e^{\int_{\delta^{\prime}T}^{s}2\lambda(\psi^{\nu}_{t})\mathrm{d}t}\leq\mathbb{E}[|Y^{\theta}_{s}|^{2}]\,,$
and that integrated over $s\in[\delta^{\prime}T,\delta T]$ reads as
@@
From Itô’s formula and (2.2), by taking expectation we have
@@
which integrated over $s\in[0,\delta^{\prime}T]$ yields to
@@
Our last ingredient is an upper bound for the time integral of the third derivative term $\mathbb{E}\|\nabla^{3}\psi^{\nu}_{t}(X^{\psi^{\mu},\rho}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}$.
[…]
@@
where $\gamma_{\tau_{u}}\coloneqq\sup_{s\in[0,{\tau_{u}}]}\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}\psi^{\nu}_{s}\|_{\mathrm{HS}}$ and
$\mathcal{I}^{\psi^{\nu}}(t,s)\coloneqq\int_{t}^{s}\exp\biggl{(}\int_{t}^{u}2\lambda(\psi^{\nu}_{l})\mathrm{d}l\biggr{)}\mathrm{d}u\,.$
@@
Fix $x,\widehat{x}\in\mathbb{R}^{d}$. Our aim is controlling with $|x-\widehat{x}|$ the following quantity
$\|\nabla^{2}\psi^{\nu}_{t}(x)-\nabla^{2}\psi^{\nu}_{t}(\widehat{x})\|_{\rm HS}\,.$
In view of this, let us consider the processes $X^{t,x}_{\cdot}$ and $X^{t,\widehat{x}}_{\cdot}$ satisfying for $s\in[t,{\tau_{u}}]$
$\begin{cases}\mathrm{d}X^{t,x}_{s}=-\nabla\psi^{\nu}_{s}(X^{t,x}_{s})\,\mathrm{d}s+\mathrm{d}B_{s}\,,,\\ \mathrm{d}X^{t,\widehat{x}}_{s}=-\nabla\psi^{\nu}_{s}(X^{t,\widehat{x}}_{s})\,\mathrm{d}s+\mathrm{d}\hat{B}_{s}\quad\forall\;t\in[0,{\tau}_{\mathrm{st}})\;\text{ and }X^{t,\widehat{x}}_{s}=X^{t,x}_{s}\quad\forall\;s\geq{\tau}_{\mathrm{st}}\\ X^{t,x}_{t}=x\text{ and }X^{t,\widehat{x}}_{t}=\widehat{x}\,,\end{cases}$
@@
$\mathrm{d}\hat{B}_{s}\coloneqq(\operatorname{Id}-2\,e_{s}\,e_{s}^{{\sf T}}\,\mathbf{1}_{\{s<{\tau}_{\mathrm{st}}\}})\,\mathrm{d}B_{s}\qquad\text{where}\quad e_{s}\coloneqq\begin{cases}\frac{X^{t,x}_{s}-X^{t,\widehat{x}}_{s}}{|X^{t,x}_{s}-X^{t,\widehat{x}}_{s}|}\quad&\text{ when }r_{t}>0\,,\\ u\quad&\text{ when }r_{t}=0\,,\end{cases}$
where $r_{t}\coloneqq|X^{t,x}_{s}-X^{t,\widehat{x}}_{s}|$ and $u\in\mathbb{R}^{d}$ is a fixed (arbitrary) unit-vector. By Lévy’s characterization, $(\hat{B}_{t})_{t\geq 0}$ is a $d$-dimensional Brownian motion, therefore $X^{t,x}_{\cdot}$ and $X^{t,\widehat{x}}_{\cdot}$ are two Schrödinger bridge processes (from $\rho$ to $\nu$) started respectively in $x$ and $\widehat{x}$, coupled via the coupling by reflection.
Let us also consider the processes $Z_{s}=\nabla^{2}\psi^{\nu}_{s}(X_{s}^{t,x})$ and $\widehat{Z}_{s}=\nabla^{2}\psi^{\nu}_{s}(X_{s}^{t,\widehat{x}})$. Since
@@
Therefore, if we set $\mathrm{d}M_{s}\coloneqq\nabla^{3}\psi^{\nu}(X^{t,x}_{s})\mathrm{d}B_{s}-\nabla^{3}\psi^{\nu}(X^{t,\widehat{x}}_{s})\mathrm{d}\hat{B}_{s}$, from Itô’s formula we may firstly deduce that
@@
where the $A\cdot B$ corresponds to the Hilbert–Schmidt scalar product between the two matrices $A,\,B$ that is the scalar $\sum_{i,j}A^{ij}B^{ij}$.
From another application of Itô’s formula [^3] Here we would apply Itô’s formula to the square root function $r(a)=\sqrt{a}$. Since there is a singularity in the origin, to be more precise we should apply a standard approximation argument by firstly applying Itô’s formula to $r_{\varepsilon}(a)\coloneqq\sqrt{a+\varepsilon}$ and then let $\varepsilon\downarrow 0$ and use the Dominated Convergence Theorem. Since we have already portrayed this approximation in the proof of Section 3.2, we omit it here. we then have
$\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}=\frac{(Z_{s}-\widehat{Z}_{s})\cdot(Z_{s}^{2}-\widehat{Z}_{s}^{2})}{\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}}\mathrm{d}s+\frac{Z_{s}-\widehat{Z}_{s}}{\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}}\cdot\mathrm{d}M_{s}$
$+\frac{\sum_{i,j}\mathrm{d}[M^{ij}_{\cdot}]_{s}}{2\,\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}}-\frac{(Z_{s}-\widehat{Z}_{s})^{2}\cdot\mathrm{d}[M_{\cdot}]_{s}}{2\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}^{3}}\,.$
@@
$(Z_{s}-\widehat{Z}_{s})\cdot(Z_{s}^{2}-\widehat{Z}_{s}^{2})=(Z_{s}-\widehat{Z}_{s})\cdot(Z_{s}-\widehat{Z}_{s})(Z_{s}+\widehat{Z}_{s})+(Z_{s}-\widehat{Z}_{s})\cdot(\widehat{Z}_{s}Z_{s}-Z_{s}\widehat{Z}_{s})$
$=(Z_{s}-\widehat{Z}_{s})^{2}\cdot(Z_{s}+\widehat{Z}_{s})\geq-\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}\,\|Z_{s}+\widehat{Z}_{s}\|_{\mathrm{HS}}\geq-2\gamma_{\tau_{u}}\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}\,.$
@@
Putting these two remarks together yields to
@@
Henceforth, the rest of the proof deals with estimating $\mathbb{P}(X^{t,x}_{s}\neq X^{t,\widehat{x}}_{s})$ for any $s\in[t,\tau_{u}]$. To do so we look at the one-dimensional process $r_{s}=|X_{s}^{t,x}-X_{s}^{t,\widehat{x}}|$, so that $\mathbb{P}(X^{t,x}_{s}\neq X^{t,\widehat{x}}_{s})=\mathbb{P}(r_{s}>0)$. From Itô’s formula we get
$\mathrm{d}r_{s}^{2}=(-2(X_{s}^{t,x}-X_{s}^{t,\widehat{x}})(\nabla\psi^{\nu}_{s}(X_{s}^{t,x})-\nabla\psi^{\nu}_{s}(X_{s}^{t,\widehat{x}}))+4)\mathrm{d}s+4\,r_{s}\,\mathrm{d}W_{s}$
where $\mathrm{d}W_{s}=e_{s}^{\intercal}\mathrm{d}B_{s}$ is a one-dimensional Brownian motion. Therefore another application of Itô’s formula yields to
@@
Therefore the process $r_{\cdot}$ is dominated from above by the process $\widetilde{r}_{\cdot}$ which solves for $s\in[t,{\tau_{u}}]$
$\mathrm{d}\widetilde{r}_{s}=-\lambda(\psi^{\nu}_{s})\,\widetilde{r}_{s}\,\mathrm{d}s+2\,\mathrm{d}W_{s}\,,\qquad\widetilde{r}_{t}=|x-\widehat{x}|\,.$
@@
$\mathrm{d}N_{s}=2\exp\biggl{(}\int_{t}^{s}\lambda(\psi^{\nu}_{u})\mathrm{d}u\biggr{)}\mathrm{d}W_{s},\quad\text{ with }N_{t}=|x-\widehat{x}|\,.$
@@
$[N]_{s}=4\int_{t}^{s}\exp\biggl{(}\int_{t}^{u}2\lambda(\psi^{\nu}_{l})\mathrm{d}l\biggr{)}\mathrm{d}u\,.$
@@
$\mathbb{P}(X_{s}^{t,x}\neq X_{s}^{t,\widehat{x}})=$ $\,\mathbb{P}(r_{s}>0)=\mathbb{P}\Big{(}\inf_{u\in[t,s]}r_{u}>0\Big{)}\leq\mathbb{P}\Big{(}\inf_{u\in[t,s]}\widetilde{r}_{u}>0\Big{)}\leq\mathbb{P}\Big{(}\inf_{u\in[t,s]}N_{u}>0\Big{)}$
$=$ $\,\mathbb{P}\Big{(}\inf_{u\in[t,s]}B_{[N]_{u}}>-|x-\widehat{x}|\Big{)}=\mathbb{P}\Big{(}\inf_{u\in[t,[N]_{s}]}B_{u}>-|x-\widehat{x}|\Big{)}$
$=$ $\,\mathbb{P}\Big{(}\sup_{u\in[t,[N]_{s}]}B_{u}\leq|x-\widehat{x}|\Big{)}=\mathbb{P}\Big{(}|B_{[N]_{s}}|\leq|x-\widehat{x}|\Big{)}$
@@
We are now ready to prove the general quantitative stability result for the Hessians. Our estimates will depend on two free parameters $\delta^{\prime}<\delta\in[0,1]$. A priori one could simply optimize over their choice; however, this optimization heavily depends on the semiconcavity parameters. We state the general result keeping these two free parameters and then choose them appropriately in Appendix B, in a setting-wise manner.
@@
with $A$ defined at (3.5).
@@
$\gamma_{\tau_{u}}\coloneqq\sup_{s\in[0,\tau_{u}]}\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}\psi^{\nu}_{s}\|_{\mathrm{HS}}\,,\quad\text{ and }\bar{\lambda}_{\tau_{\ell}}\coloneqq(\inf_{s\in[0,{\tau_{\ell}}]}\lambda(\psi^{\nu}_{s}))^{-}\,$
and recall from (3.2) (combined with (3.1)) the constant
$K_{\delta^{\prime}\delta}^{\rho\nu}=\,\frac{2\,T}{\mathcal{I}^{\psi^{\nu}}(\tau_{\ell},\tau_{u})}\,\biggl{(}3\,\Lambda(\varphi^{\nu}_{0})+\frac{\delta}{1-\delta}+2\,\sqrt{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}\biggr{)}$
+ $+\,4\,T\,\bar{\lambda}_{\tau_{\ell}}\biggl{(}3\,\Lambda(\varphi^{\nu}_{0})+\frac{\delta^{\prime}}{1-\delta^{\prime}}+2\,\sqrt{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}\biggr{)}.$
From Section 3.2 and Section 2.1 we see that

### v2

@@ 3.1. Quantitative stability estimates of Hessian
Let us consider once again the function $\theta_{s}\coloneqq\psi^{\nu}_{s}-\psi^{\mu}_{s}$ introduced in Section 2 and the forward process $(Y^{\theta}_{s})_{s\in[0,T]}$ defined as $Y^{\theta}_{s}\coloneqq\nabla\theta_{s}(X^{\psi^{\mu},\rho}_{s})$, where $(X^{\psi^{\mu},\rho})_{s\in[0,T]}$ is the Schrödinger bridge (1.4) (from $\rho$ to $\mu$), and recall that
@@
$\mathrm{d}Z^{\theta}_{s}=\Big[2\,\mathrm{sym}(Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\mu}(X^{\psi^{\mu},\rho}_{s}))+\nabla^{3}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}\Big]\,\mathrm{d}s+\nabla^{3}\theta_{s}(X^{\psi^{\mu},\rho}_{s})\mathrm{d}B_{s}\,$,
where for any square matrix $M$ the symbol $\mathrm{sym}(M)\coloneqq(M+M^{\intercal})/2$ denotes its symmetrized version and where for any $h\in\{\psi^{\mu}_{s},\,\theta_{s}\}$ and $v\in\mathbb{R}^{d}$ we have defined the product $\nabla^{3}h\,v$ as the matrix with entries $(\nabla^{3}h\,v)_{ij}:=\langle\nabla(\partial_{i}\partial_{j}h),\,v\rangle$.
@@
$2$ $\frac{Z^{\theta}_{s}\cdot\mathrm{sym}(Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s}))}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}=\frac{\operatorname{Tr}(Z^{\theta}_{s}\cdot Z^{\theta}_{s}\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s}))+\operatorname{Tr}(Z^{\theta}_{s}\cdot\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Z^{\theta}_{s})}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}$
$\,=\frac{2}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\,\operatorname{Tr}(Z^{\theta}_{s}\cdot\nabla^{2}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Z^{\theta}_{s})\geq\frac{2\,\lambda(\psi^{\nu}_{s})\,\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\geq 2\,\lambda(\psi^{\nu}_{s})\,\|Z^{\theta}_{s}\|_{\mathrm{HS}}-2\,\lambda(\psi^{\nu}_{s})\,\varepsilon\,.$
@@
$\frac{Z^{\theta}_{s}\cdot\nabla^{3}\psi_{s}^{\nu}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}=\frac{\sum_{ijk}Z^{\theta,ij}_{s}\partial_{ijk}\psi^{\nu}_{s}(X^{\psi^{\mu},\rho}_{s})Y^{\theta,k}_{s}}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\geq-\|\nabla^{3}\psi^{\nu}_{s}(X^{\psi^{\mu}}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\,.$
@@
$\mathrm{d}r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})\geq\biggl(-\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}+2\,\lambda(\psi^{\nu}_{s})\,\|Z^{\theta}_{s}\|_{\mathrm{HS}}-\|\nabla^{3}\psi^{\nu}_{s}(X^{\psi^{\mu}}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}-2\,\lambda(\psi^{\nu}_{s})\,\varepsilon\biggr)\mathrm{d}s$
$+\frac{Z^{\theta}_{s}\cdot\nabla^{3}\theta_{s}(X^{\psi^{\mu},\rho}_{s})}{r_{\varepsilon}(\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2})}\mathrm{d}B_{s}\,.$
@@
$\mathbb{E}\|Z^{\theta}_{0}\|_{\mathrm{HS}}\leq\mathbb{E}\|Z^{\theta}_{t}\|_{\mathrm{HS}}+\int_{0}^{t}\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}\mathrm{d}s-2\int_{0}^{t}\lambda(\psi^{\nu}_{s})\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}\mathrm{d}s+\int_{0}^{t}\mathbb{E}\|\nabla^{3}\psi^{\nu}_{t}(X^{\psi^{\mu},\rho}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\mathrm{d}s\,.$
@@
Applying Jensen’s inequality concludes our proof.
@@
which generalizes the constant $C^{\psi^{\nu}}_{\delta}$ considered in Theorem 3.1. By repeating the same argument employed in Theorem 2.1 when proving the upper bound (2.6b) for $\mathbb{E}[|Y^{\theta}_{T}|^{2}]=\mathbb{E}[|Y^{\eta}_{0}|^{2}]$, we can prove the following generalization.
Lemma $3.$
@@
By reasoning as in the proof of Theorem 2.1, from Itô’s formula and (2.1b), we have
$\mathrm{d}\mathbb{E}[|Y^{\theta}_{s}|^{2}]\geq 2\,\mathbb{E}[Y^{\theta}_{s}\cdot\nabla^{2}\psi^{\nu}_{s}(X^{\psi^{\mu},\rho}_{s})Y^{\theta}_{s}]\mathrm{d}s\geq 2\lambda(\psi^{\nu}_{s})\,\mathbb{E}[|Y^{\theta}_{s}|^{2}]\mathrm{d}s\,,\quad\forall s\leq\delta T\,$,
which combined with Grönwall’s lemma gives $\mathbb{E}[|Y^{\theta}_{\delta^{\prime}T}|^{2}]\,e^{\int_{\delta^{\prime}T}^{s}2\lambda(\psi^{\nu}_{t})\mathrm{d}t}\leq\mathbb{E}[|Y^{\theta}_{s}|^{2}]$.
Integrating this inequality over $s\in[\delta^{\prime}T,\delta T]$ gives
@@
From Itô’s formula and (2.1b), by taking expectation we have
@@
which integrated over $s\in[0,\delta^{\prime}T]$ leads to
@@
Our last ingredient is an upper bound for the time integral of the third derivative term.
Proposition 3.5.
@@
where
$\gamma_{\tau_{u}}\coloneqq\sup_{s\in[0,{\tau_{u}}]}\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}\psi^{\nu}_{s}\|_{\mathrm{HS}}\quad\textrm{and}\quad\mathcal{I}^{\psi^{\nu}}(t,s)\coloneqq\int_{t}^{s}\exp\biggl(\int_{t}^{u}2\lambda(\psi^{\nu}_{l})\mathrm{d}l\biggr)\mathrm{d}u\,.$
@@
Fix $x,\widehat{x}\in\mathbb{R}^{d}$. Our aim is controlling $\|\nabla^{2}\psi^{\nu}_{t}(x)-\nabla^{2}\psi^{\nu}_{t}(\widehat{x})\|_{\rm HS}$ with |xx^||x-$\widehat${x}|. In view of this, let us consider the processes $X^{t,x}_{\cdot}$ and $X^{t,\widehat{x}}_{\cdot}$ satisfying for $s\in[t,{\tau_{u}}]$
[…]
@@
[…]
@@
Therefore, if we set $\mathrm{d}M_{s}\coloneqq\nabla^{3}\psi^{\nu}(X^{t,x}_{s})\mathrm{d}B_{s}-\nabla^{3}\psi^{\nu}(X^{t,\widehat{x}}_{s})\mathrm{d}\hat{B}_{s}$, from Itô’s formula we first deduce that
@@
where the $A\cdot B$ corresponds to the Hilbert–Schmidt scalar product between the two matrices A,BA,\,B that is the scalar $\sum_{i,j}A^{ij}B^{ij}$.
From another application of Itô’s formula (as we already did in the proof of Section 3.1, by firstly applying it to $r_{\varepsilon}(a)\coloneqq\sqrt{a+\varepsilon}$ and then let $\varepsilon\downarrow 0$ ) we then have
$\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}=\frac{(Z_{s}-\widehat{Z}_{s})\cdot(Z_{s}^{2}-\widehat{Z}_{s}^{2})}{\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}}\mathrm{d}s+\frac{(Z_{s}-\widehat{Z}_{s})\cdot\mathrm{d}M_{s}}{\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}}+\frac{\sum_{i,j}\mathrm{d}[M^{ij}_{\cdot}]_{s}}{2\,\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}}-\frac{(Z_{s}-\widehat{Z}_{s})^{2}\cdot\mathrm{d}[M_{\cdot}]_{s}}{2\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}^{3}}\,.$
@@
$(Z_{s}-\widehat{Z}_{s})\cdot(Z_{s}^{2}-\widehat{Z}_{s}^{2})=(Z_{s}-\widehat{Z}_{s})^{2}\cdot(Z_{s}+\widehat{Z}_{s})\geq-\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}\,\|Z_{s}+\widehat{Z}_{s}\|_{\mathrm{HS}}\geq-2\gamma_{\tau_{u}}\|Z_{s}-\widehat{Z}_{s}\|_{\mathrm{HS}}\,.$
@@
Putting these two remarks together yields
@@
[…]
$\mathrm{d}r_{s}^{2}=(-2(X_{s}^{t,x}-X_{s}^{t,\widehat{x}})(\nabla\psi^{\nu}_{s}(X_{s}^{t,x})-\nabla\psi^{\nu}_{s}(X_{s}^{t,\widehat{x}}))+4)\mathrm{d}s+4\,r_{s}\,\mathrm{d}W_{s}\,$,
where $\mathrm{d}W_{s}=e_{s}^{\intercal}\mathrm{d}B_{s}$ is a one-dimensional Brownian motion. Therefore another application of Itô’s formula yields
@@
Hence the process $r_{\cdot}$ is dominated from above by the process $\widetilde{r}_{\cdot}$ which solves for $s\in[t,{\tau_{u}}]$
$x^|.\mathrm{d}\widetilde{r}_{s}=-\lambda(\psi^{\nu}_{s})\,\widetilde{r}_{s}\,\mathrm{d}s+2\,\mathrm{d}W_{s}\,,\qquad\widetilde{r}_{t}$=|x-$\widehat${x}|\,.
@@
$x^|.\mathrm{d}N_{s}=2\exp\biggl(\int_{t}^{s}\lambda(\psi^{\nu}_{u})\mathrm{d}u\biggr)\mathrm{d}W_{s},\quad\text{ with }N_{t}$=|x-$\widehat${x}|\,.
@@
$N]_{s}=4\int_{t}^{s}\exp\biggl(\int_{t}^{u}2\lambda(\psi^{\nu}_{l})\mathrm{d}l\biggr)\mathrm{d}u\,.$
@@
[…]
@@
We are now ready to prove the general quantitative stability result for the Hessians.
@@
with AA defined at (3.5) and $K_{\delta^{\prime}\delta}^{\rho\nu}$ defined at (3.2).
@@
$\gamma_{\tau_{u}}\coloneqq\sup_{s\in[0,\tau_{u}]}\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}\psi^{\nu}_{s}\|_{\mathrm{HS}}\,,\quad\text{ and }\bar{\lambda}_{\tau_{\ell}}\coloneqq(\inf_{s\in[0,{\tau_{\ell}}]}\lambda(\psi^{\nu}_{s}))^{-}\,.$
From Section 3.1, Section 2.1 and Hölder’s inequality we see that